The following quarterly information is available for Tons Mined Quarter Tons Mined Utility Cost Year 1-First 15,000 50,000 Year 1-Second 11,000 45,000 Year 1--Third 21,000 60,000 Year 1--Fourth 12,000 75,000 Year 2-First 18,000 100,000 Year 2-Second 25,000 105,000 Year 2--Third 30,000 $85,000 Year 2--Fourth 28,000 120,000
The following quarterly information is available for Direct Labor Hours Quarter DLH's Utility Cost Year 1-First 5,000 50,000 Year 1-Second 3,000 S 45,000 Year 1--Third 4,000 $ 60,000 Year 1--Fourth 6,000 75,000 Year 2-First 10,000 100,000 Year 2-Second 9,000 105,000 Year 2--Third 8,000 $ 85,000 Year 2--Fourth 11,000 120,000
Hard Rock is trying to estimate Utility Cost of the First quarter of Year 3. If Hard Rock uses Direct Labor Hours as the independent variable and 7,000 DLH's are anticipated, compute the amount of Utility Cost. Show computations.
Answer #1

Variable utility cost per direct labor hour = Difference in high low cost/Difference in high low activity

= (120,000 - 45,000)/(11,000 - 3,000)

= 75,000/8,000

= $9.375

Hence, total variable utility cost at 3,000 direct labor hours = 3,000 x 9.375

= $28,125

Fixed utility cost at 3,000 direct labor hours = Total utility cost at 3,000 direct labor hours - Total variable utility cost at 3,000 direct labor hours

= 45,000 - 28,125

= $16,875

Total utility cost at 7,000 direct labor hours = 16,875 + 7,000 x 9.375

= 16,875 + 65,625

= $82,500

Kindly comment if you need further assistance. Thanks